OpenStudy (anonymous):

Find an equation to the tangent line to the curve at the given point. (So, I have to find the derivative, but Im stuck - not sure if I am doing this correctly). Function is y=x^(5/2) Point is \(\ (4,32). Please show me step by step! I'm really confused!

OpenStudy (anonymous):

you can do it the usual way.... (x^n) ' = nx^(n-1) n is 5/2

OpenStudy (anonymous):

Okay, so that is the derivative....?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I thought it was 5/2(x)^(3/2)

OpenStudy (anonymous):

Where do I proceed from here?

OpenStudy (anonymous):

plug in x=4 to find the slope (m) of the tangent to the curve at that point.. use y= mx +b with the point they gave in order to find b

OpenStudy (anonymous):

Oh Okay. I was trying to plug the derivative in for m, but I guess that isn't correct.

OpenStudy (anonymous):

yeah, you just want the value of the derivative at that particular point...

OpenStudy (anonymous):

not the general expression for the slope at any value of x..

OpenStudy (anonymous):

So Is this the derivative? 5/2(x)^(3/2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks for the help!

OpenStudy (anonymous):

sure:)